OpenStudy (ofmiceparade):

Rewrite f(x) = 2(x − 1)2 + 3 from vertex form to standard form.

OpenStudy (ofmiceparade):

Hey @Learner11

OpenStudy (ofmiceparade):

I tried doing this and i can't figure it out

OpenStudy (anonymous):

hello @ofmiceparade :D

OpenStudy (ofmiceparade):

help?

OpenStudy (anonymous):

i am trying my best

OpenStudy (ofmiceparade):

f(x) = 2x2 + 5 f(x) = 4x2 − 8x + 7 f(x) = 2x2 + 4x + 5 f(x) = 2x2 − 4x + 5 are the answer choices

OpenStudy (welshfella):